OpenStudy (anonymous):

What is the quotient: (3x3 + 14x2 + 13x – 6) ÷ (3x – 1) ?

OpenStudy (anonymous):

x^2 + 14/3 x - 13/3 + 6 = x^2 + 14/3 x - 13/3 + 18/3 = x^2 + 14/3 x +5/3 ??? maybe

OpenStudy (kropot72):

The quotient = x^2 + 5x + 6 = (x + 3)(x + 2)